### v2

# Revisit What You See: Disclose Language Prior in Vision Tokens for LVLM Decoding

Abstract
Large Vision–Language Models (LVLMs) achieve strong performance across multimodal tasks by integrating visual perception with language understanding. However, how vision information contributes to the model’s decoding process remains under-explored, as reflected in frequent hallucinations. Through a series of analyses, we found that (i) vision tokens provide meaningful visual information even when hallucinations occur, and (ii) their semantics are encoded in the textual space and become explicit under appropriate vocabulary constraints. Building on these observations, we propose ReVisiT, a simple training-free decoding method that references vision tokens to guide text generation. Our approach leverages the semantic information embedded within vision tokens by projecting them into the text token distribution. Specifically, ReVisiT dynamically selects the most relevant vision token at each decoding step via context-aware constrained divergence minimization, and using its constrained projection to refine the output distribution to better incorporate visual semantics. Across five benchmarks on recent LVLMs, ReVisiT consistently enhances visual grounding with minimal computational overhead, and achieves competitive or superior results to state-of-the-art decoding baselines while reducing computational cost by up to $2\times$.  [^1] Code is available at https://github.com/bscho333/ReVisiT.
1 Introduction
With the recent success of Large Language Models (LLMs) (Touvron et al., 2023; Achiam et al., 2023; Team et al., 2023), Large Vision-Language Models (LVLMs) have emerged as powerful multimodal architectures that integrate visual perception with language understanding and text generation (Dai et al., 2023; Liu et al., 2024a; Zhu et al., 2024; Bai et al., 2025; Zhu et al., 2025). Typically, LVLMs encode visual inputs into the LLM decoder’s text embedding space as vision tokens, which are then processed alongside text tokens during decoding. This way of treating vision tokens as static auxiliary contexts, similar to retrieval-augmented generation (Lewis et al., 2020; Gao et al., 2023) in LLMs, enables the construction of complex multimodal systems in a relatively simple manner. However, this approach is often insufficient to capture the unique characteristic and role of vision tokens as sole carriers of visual information, as denoted by hallucinations frequently observed in LVLM’s text outputs (Leng et al., 2024; Favero et al., 2024; Huo et al., 2025; Rohrbach et al., 2018; Li et al., 2023b; Guan et al., 2024). To mitigate this, research has begun to expand the understanding of vision tokens along with their additional utilization (Jiang et al., 2025a; b). Yet, it is still under-explored how vision tokens influence the decoding process of LVLMs and what kinds of textual semantics they encode.
To investigate this direction, we conduct a series of analyses on vision tokens, leading to two key insights. First, by examining the output distributions of LVLMs particularly at hallucinated steps, we find that ground-truth objects often remain among the model’s high-probability candidates and this visual grounding signal in output distribution indeed comes from the vision tokens. Second, we interpret the semantics of vision tokens by projecting them into the text token distribution using the LLM decoder’s language modeling head, as both are embedded in the same space. Here, we observe that the textual semantics of vision tokens is initially obscured, but the interpretable and well-aligned semantics could be discovered in the projected distribution with the proper constraints to focus on specific subset. Overall, these findings suggest that vision tokens intrinsically encode object-level semantics, though such signals typically remain latent under conventional unconstrained decoding.
Building on this observation, we propose ReVisiT, a simple yet effective decoding strategy that Referencing Vision Tokens to guide the text generation process of LVLMs. At each step, ReVisiT selects the most relevant vision token based on context-aware constrained divergence from the current output distribution, and uses its constrained projection as reference to refine token-level logits. This activates the intrinsic semantics of vision tokens without requiring additional training, architectural modifications, external modules, or multi-pass inference.
Figure 1: ReVisiT across various model sizes and architectures. We evaluate on the CHAIR benchmark and report F1. Consistent improvements hold across different size buckets (7–8B / 13–14B / 26–32B) and architectures, demonstrating scalability and model-agnostic effectiveness. Full results are provided in Table 9.
As highlighted in Fig. 1,  ReVisiT consistently improves performance across six model sizes spanning three architecture, demonstrating scalability and model-agnostic effectiveness. To be specific, on three hallucination benchmarks—HallusionBench (Guan et al., 2024), CHAIR (Rohrbach et al., 2018), and POPE (Li et al., 2023b)—ReVisiT reduces hallucinations robustly on both LLaVA-1.5 (Liu et al., 2024a) and Qwen2.5-VL (Bai et al., 2025). Beyond hallucination, we further validate its effectiveness on general-purpose benchmarks, including VQAv2 (Antol et al., 2015) and MMMU (Yue et al., 2024), indicating that the semantics surfaced by our context-aware constrained projection generalize across tasks and architectures.
In summary, our contributions are as follows:
  • $\circ$
    We provide quantitative and qualitative evidences that vision tokens intrinsically encode interpretable semantics exposed when projected over a semantically coherent vocabulary subset.
  • $\circ$
    We introduce ReVisiT, a simple yet effective LVLM decoding method that Referencing Vision Tokens via context-aware constrained divergence.
  • $\circ$
    We validate ReVisiT across six model sizes spanning three architectures, reducing hallucinations and yielding gains on five benchmarks, while maintaining inference cost.
2 Motivation
Figure 2: Motivation. We qualitatively analyzed various vision tokens with LLaVA-1.5-7B (Liu et al., 2024a). Dotted arrows represent vision token projection over specified vocabulary set. For each box, representing text token distribution, we annotated top-5 probable text tokens. Left part illustrate the effectiveness of vocabulary constraint, whereas right part shows the distribution shift during ReVisiT. See Appendix C.1 for a detailed discussion of the underlying values and analysis.
In this section, we investigate whether vision tokens encode interpretable semantics in LVLM’s text token space that can directly inform its decoding process. To conduct systematic analyses, we adopt the CHAIR benchmark (Rohrbach et al., 2018), as it provides a principled definition of hallucinations based on 80 MSCOCO (Lin et al., 2014) object categories with 403 synonyms in total; for each sample, the set of synonyms of annotated object categories is regarded as ground-truth (GT) objects, while the remaining category synonyms are regarded as hallucinated (Hal) objects. We refer to a hallucinated step as a decoding step where generated token corresponds to an object in this MSCOCO synonym set but is not part of the ground-truth annotations for the given image.
Building on this definition, we sample 500 images from the MSCOCO validation set and generate captions for all images. All quantitative analyses are performed with Qwen2.5-VL-7B (Bai et al., 2025) using vanilla greedy decoding with a maximum of 512 tokens, prompted with "Please describe this image in detail.", which is a standard input prompt for CHAIR.
2.1 How do vision tokens affect the output distribution of LVLM?
To examine how vision tokens affect the decoding step of LVLMs, we first ask whether ground-truth (GT) objects remain among the model’s top-probability predictions at hallucinated steps. We use a at-least-one recall@k defined on a probability distribution over the vocabulary: for a given distribution, we measure whether at least one GT object synonym appears within its top-$k$ predictions. Among the 500 generated captions, 176 captions contain at least one hallucinated step, and we obtain 190 hallucinated steps for step-level analysis. For each hallucinated step, we report at-least-one recall@k on the model’s output distribution at that step, excluding the greedy (top-1) prediction from the candidate set (details in Appendix A.1.2).
Figure 3: Ground-truth objects frequently remain in top-probability predictions. At 190 hallucinated steps, at least one GT object is recalled in 95.8% of cases within the top-500 predictions, corresponding to only 0.33% of full vocabulary (151,665 tokens).
The results are summarized in Fig. 3. It is observable that GT objects frequently remain accessible in output distribution: in 63.2% of hallucinated steps, at least one GT object appears within the top-50 candidates, and this recall rises to 95.8% within the top-500. Given the vocabulary size of 151,665 tokens, top-500 corresponds to only 0.33% of the full vocabulary. A qualitative example is also presented in Fig. 2 (right), where the model incorrectly selects “person” as the greedy prediction, yet visually grounded alternatives such as “mountain,” “landscape,” and “painting” appear among the top-5 candidates. This indicates that relevant visual candidates are assigned high probabilities in the output distribution, showing that the LVLM indeed possesses the correct visual information even when hallucinations occur. The 95% confidence intervals are reported in Appendix A.1.2.
We further conduct a fine-grained comparison between the occurrences of GT and hallucinated object subsets during the decoding. Specifically, for each generated token, we track the maximum probability among objects in each subset. Then, we report the average of these per-step maxima as GT Max Avg for GT object subsets and Hal Max Avg for hallucinated object subsets, respectively (see detailed definition in Appendix A.1.3). To investigate the contribution of vision tokens by isolating them, we additionally report w/o image as an image-absent setting where we remove all vision tokens and perform teacher forcing with the previously generated caption as input while keeping decoding otherwise identical.
Table 1: Comparison of Hal Max Avg and GT Max Avg. The results for (i) all output tokens across all 500 CHAIR samples (88,706 tokens), (ii) all output tokens within hallucinated samples (176 samples, 35,116 tokens), and (iii) hallucinated decoding steps within hallucinated samples (176 samples, 190 steps). The higher value between Hal and GT is highlighted in bold.
Analysis Setting Condition GT Max Avg Hal Max Avg
All tokens w/ image 24.49 20.24
w/o image 21.83 22.66
All tokens w/ image 24.42 21.74
(in hallucinated samples) w/o image 21.73 23.88
Hallucinated tokens w/ image 32.00 36.44
w/o image 30.26 35.23
As shown in Table 1, with vision tokens present, GT objects receive higher probability mass than hallucinated ones on average (GT >> Hal); removing the image tokens reverses the ordering (Hal >> GT), confirming a grounding contribution from visual input. At hallucinated steps, the reversed ordering (Hal >> GT) persists even with the presentation of image, indicating that language priors (i.e., textual co-occurrence biases from pre-training (Leng et al., 2024; Favero et al., 2024)) can dominate the prediction at specific positions. We further validate this ordering shift using a paired nonparametric Wilcoxon signed-rank test (Wilcoxon, 1945), which compares the signs and magnitudes of paired step-level differences without assuming normality, showing a significant shift toward GT when moving from w/o-image to w/-image condition in the full-token analyses (Table 1, settings (i) and (ii)), while no significant shift is observed on the hallucinated-step subset (See Appendix A.1.3 for detailed discussion).
2.2 What textual semantics are encoded in vision tokens?
To identify the source of visual grounding signals, we focus on vision tokens, the sole carriers of visual information in LVLMs. Vision tokens share the same embedding space with text tokens and are autoregressively processed during decoding like them. This motivates probing their representations by projecting them through the model’s language-modeling head to obtain the text token distribution over the vocabulary, analogous to logit-lens and early-exit analyses for text token embeddings (nostalgebraist, 2020; Chuang et al., 2024; Wang et al., 2025). Specifically, for each of 500 MSCOCO validation images, we project each vision token’s final-layer hidden state to obtain the text token distribution, yielding 173,173 vision token projections. (see details in Appendix A.1.4).
Figure 4: Vision token projection. Vision tokens reveal rich semantics when projected over semantically coherent subsets.
First, to assess the presence of visual grounding signals, we compute token-level at-least-one recall@k for each vision token projection, following the definition in Sec. 2.1. Quantitatively, across 173,173 vision tokens, at-least-one recall@k is low under the full vocabulary: only 2.03% of projections surface a GT synonym at Top-1, increasing to 21.72% at Top-30. Qualitative inspection also indicates that tokens corresponding to background regions often yield generic function words, with object labels rarely appearing (red boxes on the left side of Fig. 2). These observations indicate that, under the full vocabulary, vision-token projections are diffuse and difficult to interpret.
In contrast, constraining the projection to semantically coherent vocabulary subsets reveals substantially clearer alignment with visual semantics. When restricting the prediction space to the CHAIR-derived object subset (176 single-token synonyms covering 62 categories), at-least-one recall@k on the same 173,173 projections increases sharply to 40.44% at Top-1—about $20\times$ higher than the naïve full-vocabulary projection (2.03%). At Top-30, recall reaches 89.24% while inspecting only $30/176\!\approx\!17\%$ of the subset. For comparability, both settings evaluate GT synonyms from the same 62 categories. Moreover, when we manually define domain-specific subsets, such as environment-related terms including “river,” “mountain,” and “tree”, the projected distribution aligns with visually salient regions and reveals object labels that remain hidden under the full-vocabulary setting (yellow boxes on the left side of Fig. 2).
These findings provide both quantitative and qualitative evidence that vision tokens intrinsically encode object-level semantics, although such signals are not easily observable without appropriate semantic constraints. Details are presented in Appendix A.1.4.
3 ReVisiT: Referencing Vision Tokens to Guide Text Generation
Based on insights from Sec. 2, we propose ReVisiT, a decoding method that adaptively leverages intermediate vision token representations as internal reference signals to guide text generation, enhancing visual grounding with minimal computational overhead. At each decoding step, ReVisiT focuses the decoding process onto a context-relevant vocabulary subset and refines the output by referencing the most relevant vision token. To be specific, ReVisiT consists of the following three main steps: (1) adaptively constraining the vocabulary based on the vanilla output distribution, (2) projecting and selecting vision token hidden states within the subset, and (3) refining the output distribution through element-wise combination and normalization to generate the next token.
Preliminary. Let denote input text tokens (e.g., query) and vision tokens as $\mathbf{x}$ and $\mathbf{v}$, respectively. Specifically, the vision tokens $\mathbf{v}$ are generated from the given input image, by forwarding it into vision encoder and cross-modality projector sequentially (Liu et al., 2024a; Bai et al., 2025). Then, our goal is to generate output text tokens $\mathbf{y}$ using LVLM $\mathcal{M}$; following the auto-regressive nature of LLM, output tokens $\mathbf{y}$ (e.g., response) are sequentially generated, i.e., $y_{t}\sim\mathcal{M}(\cdot|\mathbf{v},\mathbf{x},\mathbf{y}_{<t-1})$. Let assume that LVLM $\mathcal{M}$ has $L$ decoding layers and text token vocabulary $\mathcal{V}$, and denote its hidden state of $l$-th layer as $h^{l}$. The output distribution is parameterized by model weights $\theta$ and denoted as $p_{\theta}$. Then, at time step $t$, the output probability over $\mathcal{V}$ is calculated as below:
$p_{\theta}(h_{T+t-1}^{L},\mathcal{V})=\mathrm{softmax}_{\mathcal{V}}\left(\phi(h_{T+t-1}^{L})\right),~\sum_{w\in\mathcal{V}}p_{\theta}(w|h_{T+t-1}^{L},\mathcal{V})=1,$ (1)
where $T=|\mathbf{v}|+|\mathbf{x}|$, $\phi$ is the language modeling head which maps final hidden state of LVLM into logits over vocabulary $\mathcal{V}$, and $\mathrm{softmax}_{\mathcal{V}}$ indicates a softmax operation over $\mathcal{V}$. Then, the output token $y_{t}$ is sampled from the obtained distribution:
$y_{t}\sim p_{\theta}(h_{T+t-1}^{L},\mathcal{V}).$ (2)
Figure 5: An overview of ReVisiT. At each decoding step, ReVisiT (1) constrains the vocabulary $\mathcal{V}$ to $\mathcal{V}_{\texttt{cons}}^{t}$, (2) projects vision token embeddings into $\mathcal{V}_{\texttt{cons}}^{t}$ and selects most relevant token, and (3) refines the final output distribution. ReVisiT leverages vision tokens to serve as reference signals for decoding, enhancing visual grounding without additional forward passes or external supervision.
3.1 Adaptive context-aware vocabulary subset construction
At each decoding timestep $t$, LVLM $\mathcal{M}$ generate output distributions $p_{\theta}(h_{T+t-1}^{L},\mathcal{V})$ over the entire vocabulary space $\mathcal{V}$. However, as observed in Figure 2, the original distribution is often diffusely spread across many irrelevant tokens. Therefore, we dynamically constrain a vocabulary $\mathcal{V}$ into subset $\mathcal{V}_{\texttt{cons}}^{t}$ more suitable to capture the plausible semantic space of the output distribution. By restricting the output space to a smaller, contextually focused subset, this constraint reduce semantic distraction and facilitate more precise integration of visual information. To be specific, following Li et al. (2023a), we adaptively define the vocabulary subset $\mathcal{V}_{\texttt{cons}}^{t}$ as below:
$\mathcal{V}_{\texttt{cons}}^{t}=\left\{w\in\mathcal{V}:p_{\theta}(w|h_{T+t-1}^{L},\mathcal{V})\geq\alpha\cdot\max_{w^{\prime}}p_{\theta}(w^{\prime}|h_{T+t-1}^{L},\mathcal{V})\right\},$ (3)
where $\alpha\in(0,1)$ controls the sparsity of the selected candidates. This vocabulary subset $\mathcal{V}_{\texttt{cons}}^{t}$ serves as a dynamically adapted output space, which all subsequent vision token projections and refinements are performed to enhance visual grounding without introducing additional computation.
3.2 Vision token projection and selection over subset
Given the dynamically constrained vocabulary subset $\mathcal{V}_{\texttt{cons}}^{t}$ at each decoding timestep $t$, we next describe how vision token hidden states $h_{i}^{j}$ are projected and selected within this constrained space. Let $\mathcal{J}$ as the set of candidate decoder layers where vision token hidden states are selected. Specifically, for each $h_{i}^{j}$, where $i$ denotes the index of the vision tokens and $j$ denotes the candidate layer, we obtain the projected distribution over the constrained vocabulary $\mathcal{V}_{\texttt{cons}}^{t}$:
$p_{\theta}(h_{i}^{j},\mathcal{V}_{\texttt{cons}}^{t}),\quad\text{for}\quad i\in\{0,\ldots,|\mathbf{v}|-1\},\;j\in\mathcal{J}.$ (4)
Intuitively, this distribution represents the likelihood of each candidate token within $\mathcal{V}_{\texttt{cons}}^{t}$ conditioned on the vision token embedding $h_{i}^{j}$. Next, to identify the vision token most relevant to the current decoding context, we compare each vision token distribution $p_{\theta}(h_{i}^{j},\mathcal{V}_{\texttt{cons}}^{t})$ against the vanilla output distribution $p_{\theta}(h_{T+t-1}^{L},\mathcal{V}_{\texttt{cons}}^{t})$ by computing the Jensen-Shannon Divergence (JSD):
$(i^{*},j^{*})=\arg\underset{i,j}{\min}\;\mathrm{JSD}\left(p_{\theta}(h_{T+t-1}^{L},\mathcal{V}_{\texttt{cons}}^{t})\parallel p_{\theta}(h_{i}^{j},\mathcal{V}_{\texttt{cons}}^{t})\right).$ (5)
Here, $i^{*}$ and $j^{*}$ denote the indices of the selected vision token and decoder layer, respectively. We adopt JSD as the divergence metric due to its symmetric properties and bounded output range.
Remark. For the efficient implementation, the vision token projections over the full vocabulary $\mathcal{V}$, $p_{\theta}(h_{i}^{j},\mathcal{V})$ , are precomputed once before decoding begins. Then, at each timestep $t$, we apply slicing and masking to adapt the cached projections to the dynamically constrained subset $\mathcal{V}_{\texttt{cons}}^{t}$, enabling efficient selection without additional forward computation.
3.3 Output logit refinement using distribution from selected vision token
After identifying the most contextually relevant vision token at each decoding step, we then refine the vanilla output distribution guided by the reference visual grounding signal from the selected vision token. Specifically, given the vanilla output distribution $p_{\theta}(h_{T+t-1}^{L},\mathcal{V}_{\texttt{cons}}^{t})$ and the selected vision token distribution $p_{\theta}(h_{i^{*}}^{j^{*}},\mathcal{V}_{\texttt{cons}}^{t})$, we combine them through element-wise multiplication:
$p_{\tt{fin}}(y_{t})\propto\begin{cases}p_{\theta}(y_{t}|h_{T+t-1}^{L},\mathcal{V}_{\texttt{cons}}^{t})\times p_{\theta}(y_{t}|h_{i^{*}}^{j^{*}},\mathcal{V}_{\texttt{cons}}^{t}),&\text{if }y_{t}\in\mathcal{V}_{\texttt{cons}}^{t},\\ 0,&\text{otherwise}.\end{cases}$ (6)
The combined distribution $p_{\tt{fin}}$ is subsequently normalized over $\mathcal{V}_{\texttt{cons}}^{t}$ to form a valid probability distribution, and the next output token $y_{t}$ is sampled from the normalized $p_{\tt{fin}}$. This refinement mechanism allows the decoding process to dynamically integrate visual grounding signals extracted from vision tokens, without introducing additional inference passes or architectural modifications. The overall algorithm of ReVisiT is presented in Algorithm 1.
4 Experiments
4.1 Setups
Models and baselines. To validate the generality of our method across different architectures and capabilities, we conduct experiments using two representative LVLMs: LLaVA-1.5-7B (Liu et al., 2024a) as a widely adopted conventional baseline, and Qwen2.5-VL-7B (Bai et al., 2025) as a representative of state-of-the-art. As baselines, we adopt various intra-alignment decoding strategies to compare the effectiveness of our method in enhancing visual grounding. (a) Greedy decoding is a standard decoding baseline, and (b) DoLa (Chuang et al., 2024) improves factuality by contrasting output logits from earlier and later decoder layers at each timestep, and has been widely adopted for hallucination mitigation in LVLMs despite being originally proposed for LLMs. For intra-alignment baselines, (c) VCD (Leng et al., 2024) mitigates hallucination by perturbing visual inputs with Gaussian noise and enforcing consistency between the original and perturbed outputs; (d) M3ID (Favero et al., 2024) contrasts outputs with and without visual input to alleviate language prior bias and enhance visual grounding; (e) CODE (Kim et al., 2024) leverages the model’s self-generated image description as a contrasting reference to alleviate hallucination; (f) SID (Huo et al., 2025) applies contrastive decoding on low-attention vision patches to refine visual grounding.
Implementation details. In entire experiments, we apply deterministic greedy decoding (i.e., $\text{temperature}=0$, $\text{top-}p=1$, $\text{top-}k=1$, $\text{num\_beams}=1$) with a maximum output length of 512 tokens for all decoding methods. Regarding baselines, we follow the default hyperparameters provided by their original implementations. For ReVisiT, we set the hyperparameters as follows: for LLaVA-1.5-7B, we select vision tokens mostly from the last decoder layer (”last”). For Qwen2.5-VL-7B, we select vision tokens mostly from all even-numbered decoder layers (”all”). The APC threshold $\alpha$ is set for each task. More details are presented in Appendix A.4.
Table 2: HallusionBench. We report accuracies (%) on five metrics: qAcc (Question-pair accuracy), fAcc (Figure accuracy), Easy aAcc, Hard aAcc, and aAcc (All accuracy). Higher scores ($\uparrow$) indicate better performance for all metrics. he best results in each setting are bolded, and the second-best are underlined
Method LLaVA-1.5-7B Qwen2.5-VL-7B
qAcc $\uparrow$ fAcc $\uparrow$ Easy aAcc $\uparrow$ Hard aAcc $\uparrow$ aAcc $\uparrow$ qAcc $\uparrow$ fAcc $\uparrow$ Easy aAcc $\uparrow$ Hard aAcc $\uparrow$ aAcc $\uparrow$
Greedy 11.55 22.61 44.40 48.54 48.22 29.96 34.78 65.70 53.65 61.08
DoLA 13.36 22.61 45.85 49.27 49.24 27.08 30.87 62.09 51.46 58.04
VCD 14.44 22.61 44.77 47.45 47.55 29.60 35.22 64.62 54.74 61.25
M3ID 11.91 23.04 43.68 49.27 48.22 30.32 33.48 62.82 54.38 60.07
CODE 10.47 20.43 41.52 47.81 46.36 28.88 31.74 57.76 54.74 57.53
SID 12.27 21.30 44.77 48.18 48.22 32.13 36.09 65.70 55.11 61.59
[…]
Table 3: CHAIR benchmark. Lower scores ($\downarrow$) on CHAIRS, CHAIRI and higher ($\uparrow$) F1 indicate better performance. The best results in each setting are bolded, and the second-best are underlined
Method LLaVA-1.5-7B Qwen2.5-VL-7B
CHAIRS $\downarrow$ CHAIRI $\downarrow$ F1 $\uparrow$ CHAIRS $\downarrow$ CHAIRI $\downarrow$ F1 $\uparrow$
Greedy 53.8 14.66 82.33 35.2 8.43 79.85
DoLA 53.6 14.45 82.27 31.0 7.63 79.33
VCD 52.8 15.61 81.84 35.0 9.07 80.10
M3ID 57.0 16.57 82.06 33.8 9.80 78.80
CODE 45.2 13.13 81.48 17.0 9.35 48.50
SID 50.6 13.58 83.34 37.6 8.84 80.40
[…]
Table 4: POPE benchmark. Average for 3 datasets and 3 types. Higher scores ($\uparrow$) on Accuracy, Precision, and F1 indicate better performance. The best results in each setting are bolded, and the second-best are underlined. Full results are presented in Table 11.
Method LLaVA-1.5-7B Qwen2.5-VL-7B
Accuracy $\uparrow$ Precision $\uparrow$ F1 $\uparrow$ Accuracy $\uparrow$ Precision $\uparrow$ F1 $\uparrow$
Greedy 79.47 74.48 82.36 84.23 93.01 82.53
DoLA 79.58 74.67 82.41 81.36 92.36 78.71
VCD 77.74 72.51 80.97 84.43 92.48 82.91
M3ID 79.48 74.48 82.37 83.56 91.52 82.00
CODE 78.58 73.98 81.35 83.55 90.20 82.26
SID 80.40 75.94 82.91 85.78 92.27 84.72
[…]
Table 5: MMMU and VQAv2. For MMMU, we report Accuracy and Subject Win/Tie vs. Greedy; for VQAv2, we report Accuracy. Higher scores ($\uparrow$) indicate better performance for all metrics. The best results in each setting are bolded, and the second-best are underlined
Method MMMU VQAv2
LLaVA-1.5-7B Qwen2.5-VL-7B LLaVA-1.5-7B Qwen2.5-VL-7B
Accuracy $\uparrow$ Subj. Win/Tie Accuracy $\uparrow$ Subj. Win/Tie Accuracy $\uparrow$ Accuracy $\uparrow$
vs Greedy $\uparrow$ vs Greedy $\uparrow$
Greedy 34.39 - 51.11 - 69.47 65.80
DoLA 34.29 28 50.15 18 69.87 60.60
VCD 33.51 19 51.13 21 65.93 66.07
M3ID 34.51 29 46.63 15 70.33 59.87
CODE 34.48 22 42.79 5 66.00 46.60
SID 33.10 16 49.75 16 70.33 67.13
[…]
4.2 Main results
We evaluate both generative (caption faithfulness) and discriminative (object presence) grounding across five benchmarks to assess robustness and generalizability.
HallusionBench (Guan et al., 2024) probes image–context reasoning and entangled hallucinations with a LLM-based judge. We focus on the visual-dependent setting. Results are presented in Table 2, Across both LLaVA-1.5-7B and Qwen2.5-VL-7B, ReVisiT outperforms all baselines on the three representative metrics (qAcc, fAcc, and aAcc) among all methods. Notably, ReVisiT exceeds the second-best method by more than 5 points on qAcc, and reaches $1.75\times$ the score of vanilla greedy (from $11.55$ to $20.22$).
CHAIR (Rohrbach et al., 2018) measures hallucination rates in image captioning by comparing generated captions to ground-truth object annotations from MS COCO (Lin et al., 2014) dataset. We report sentence-level (CHAIRS) and instance-level (CHAIRI) hallucination scores on 500 sampled images, and additionally include F1 as an per-sample aggregate caption-level score. Table 3 shows that ReVisiT reduces hallucination while preserving or improving coverage on both models, maintaining at least second-best for all metrics.
POPE (Li et al., 2023b) evaluates object presence with binary questions across three datasets (MS COCO, A-OKVQA (Schwenk et al., 2022), and GQA (Hudson & Manning, 2019)) and three query types (random, popular, adversarial). We report Accuracy, Precision, and F1. As summarized in Table 4, ReVisiT ranks first or second in the overall averages across all datasets and query types, indicating robust gains in discriminative grounding.
MMMU (Yue et al., 2024) provides 30 subjects of VQA including both multiple-choice and open-ended settings. We report overall Accuracy and a per-subject win/tie vs. greedy rate to capture subject-wise gains over vanilla decoding. Results are presented in Table 5 (left). Across 30 subjects, ReVisiT attains the highest overall accuracy on both models while improving subject-wise robustness.
VQAv2 (Goyal et al., 2017) measures VQA correctness under inter-annotator variability via consensus accuracy over 10 annotations. As shown in Table 5 (right), ReVisiT improves general VQA performance and attains the best accuracy on both models: LLaVA-1.5-7B ($69.47\rightarrow\mathbf{70.40}$), Qwen2.5-VL-7B ($65.80\rightarrow\mathbf{67.60}$), surpassing the strongest baselines (e.g., $70.33$ and $67.13$).
Overall. ReVisiT succeeds across generative and discriminative settings, under both manually designed and LLM-aided evaluations, and on both hallucination-focused and general VQA tasks—without training, architectural changes, external modules, or multi-pass inference—by referencing intrinsic vision-token semantics. Comprehensive per-benchmark breakdowns, metrics, and extended tables are presented in Appendix A.2 and B.
4.3 Additional analyses
Inference speed improvement.
To evaluate the inference efficiency of ReVisiT compared to baseline decoding strategies, we measure the per-token computational time. All measurements are conducted on an NVIDIA RTX A6000 GPU with 48GB memory and an Intel(R) Xeon(R) Gold 5320 CPU running at 2.20 GHz. We perform 6 warmup iterations before measurement to eliminate initial overhead, and each of the 300 samples is decoded 3 times per sample during evaluation.
Figure 6: Inference speed. Comparison of per-token inference latency across different decoding strategies for LLaVA-1.5-7B (left y-axis) and Qwen2.5-VL-7B (right y-axis), with standard deviations visualized as error bars.
We report the average decoding time per generated token, measured in milliseconds (ms/token). Other evaluation configurations, including preprocessing pipelines and evaluation scripts, follow the official CHAIR benchmark settings. Exact numerical values are provided in Table 12.
As shown in Figure 6, ReVisiT matches the inference efficiency of vanilla greedy decoding, increasing time by only $1.9\%$ on LLaVA-1.5-7B and $0.6\%$ on Qwen2.5-VL-7B. By contrast, methods that require additional forward passes (e.g., VCD / M3ID / SID / CODE) approach or exceed $2\times$ the greedy runtime. Even DoLa,which only adds a projection on intermediate layers each step, incurs $29.34\%$ and $14.64\%$ overhead on LLaVA-1.5-7B and Qwen2.5-VL-7B, respectively. These results highlight the key advantage of ReVisiT: it enhances visual grounding with negligible computational overhead, unlike prior methods that significantly compromise inference efficiency.
Figure 7: Ablation study. We evaluate the impact of vocabulary subset constraint, vision token selection, layer scope, and threshold $\alpha$ variation on CHAIR benchmark, reporting F1. Comprehensive results are presented in Table 10.
Ablation study.
We ablate four factors on Qwen2.5-VL under the CHAIR benchmark (Figure 7): (1) vision-token selection criterion (min-/random/max-JSD), (2) vocabulary constraint (subset vs. full), (3) layer scope (all layers vs. last layer), and (4) APC threshold $\alpha$.
Selection. Under a constrained (subset) vocabulary, max-JSD collapses (F1 =0.67=0.67) and random underperforms (F1 =17.63=17.63), while our min-JSD attains the best performance (F1 =81.16=81.16), substantiating the effectiveness of a context-aware constrained divergence for token selection (Table 10(a)).
Vocabulary constraint. Even with the same min-JSD selection, projecting over the full vocabulary fails (F1 =1.52=1.52), exhibiting a degenerate behavior where vision-token projections are referenced off-target. This confirms that constraining the candidate set to context-relevant tokens is crucial (Table 10(a)); qualitative failures are provided in Appendix C.2.
Layer scope & threshold. Across $\alpha$, the all-layers variant peaks at $\alpha=10^{-5}$ (F1 =81.16=81.16), while the last-layer variant peaks at $\alpha=10^{-3}$ (F1 =80.97=80.97). Overall sensitivity to $\alpha$ is modest (variations within $\approx\!1.5$ F1 around each peak), indicating that all-layers + constrained vocabulary + min-JSD is a robust setting (Table 10b).
5 Conclusion
In this paper, we revisited the role of vision tokens in LVLM decoding and provided quantitative evidence that they encode interpretable semantics that become explicit under semantically coherent vocabulary constraints. Building on these findings, we introduced ReVisiT, a training-free, model-agnostic decoding strategy that references the most relevant vision token via a context-aware constrained divergence and refines token logits accordingly. As highlighted in Fig. 1, ReVisiT consistently improves performance across six model sizes spanning three architectures, demonstrating scalability and model-agnostic effectiveness.
Appendix A Detailed Discussions
A.1 Motivation
A.1.1 Metric formulation (at-least-one recall@k)
Given any probability distribution $p_{\theta}(h,\mathcal{V})$ over the vocabulary $\mathcal{V}$,
let $\mathrm{TopK}_{k}(p_{\theta})\subset\mathcal{V}$ denote the set of its top-$k$ tokens by probability, and let $G\subseteq\mathcal{V}$ be the set of ground-truth (GT) object synonyms relevant to the evaluation unit.
We define the success indicator
$I(k)\;=\;\mathbf{1}\!\left\{\,\mathrm{TopK}_{k}(q)\;\cap\;G\;\neq\;\varnothing\,\right\},$ (7)
and report the mean of $I(k)$ over all evaluation units in the set $\mathcal{U}$:
$\widehat{\mathrm{Recall}}(k)\;=\;\frac{1}{|\mathcal{U}|}\sum_{u\in\mathcal{U}}I_{u}(k).$ (8)
A.1.2 Figure 3
CHAIR Rohrbach et al. (2018) identifies 275 hallucinated words across 500 captions. Among these, we focus on the subset where the hallucinated word and at least one ground-truth (GT) synonym are both representable as single tokens by the tokenizer, resulting in $n=190$ steps available for step-level analysis. Table 6 reports the recall rates with 95% Wilson score intervals.
Table 6: At-least-one GT recall at hallucinated steps ($n=190$) with 95% Wilson intervals.
Top-$k$ % of vocab at-least-one recall@k [95% CI]
50 0.033% 63.2% [56.0, 70.0]
100 0.066% 73.7% [66.9, 79.6]
500 0.33% 95.8% [91.6, 98.1]
A.1.3 Table 1
Setup formulation.
We follow the notation in Sec. 3: at decoding step $t$, the LVLM outputs $p_{\theta}(h_{T+t-1}^{L},\mathcal{V})$ as in Eq. equation 1. For CHAIR, let $\Omega$ denote the MSCOCO synonym set ($|\Omega|=403$). To ensure the clarity of object representation and avoid ambiguity from multi-token variants, we construct the single-token subset
$\Omega^{(1)}\;=\;\bigl\{\,w\in\Omega\;:\;\texttt{tokenize}(\texttt{"\textvisiblespace{}}w\texttt{") has length }1\,\bigr\},$ (9)
which contains 176 word types covering 62 object categories. At step $t$, we define the CHAIR-based ground-truth and hallucinated synonym subsets over $\Omega^{(1)}$ as $\mathcal{G}_{t}\subseteq\Omega^{(1)}$ and $\mathcal{H}_{t}=\Omega^{(1)}\setminus\mathcal{G}_{t}$. For step-level analyses, we further require that (i) the hallucinated word at $t$ is single-tokenizable and (ii) $\mathcal{G}_{t}\neq\varnothing$, yielding $n=190$ paired steps (selection details match Appendix A.1.2). We evaluate two conditions: w/ image (vision tokens present) and w/o image (all vision tokens removed); in the latter, steps are aligned by teacher forcing with the previously generated caption while keeping all other decoding details identical.
Statistical test and reporting.
For paired step-level comparisons, we compute the maxima
$m_{t}^{\mathrm{GT}}=\max_{w\in\mathcal{G}_{t}}p_{\theta}(w\mid h_{T+t-1}^{L},\mathcal{V}),\qquad m_{t}^{\mathrm{Hal}}=\max_{w\in\mathcal{H}_{t}}p_{\theta}(w\mid h_{T+t-1}^{L},\mathcal{V}),$
and form
$d_{t}^{\mathrm{with}}=m_{t}^{\mathrm{GT}}-m_{t}^{\mathrm{Hal}},\qquad d_{t}^{\mathrm{w/o}}=m_{t}^{\mathrm{GT,w/o}}-m_{t}^{\mathrm{Hal,w/o}},\qquad\Delta_{t}=d_{t}^{\mathrm{with}}-d_{t}^{\mathrm{w/o}}.$
We apply the Wilcoxon signed-rank test (two-sided; Pratt) to $d_{t}^{\mathrm{with}}$, $d_{t}^{\mathrm{w/o}}$, and $\Delta_{t}$ on the intersection of steps available in both conditions. We report the paired sample size $N$, the median differences, and Holm-adjusted $p$-values for $\Delta_{t}$.
Table 7: Wilcoxon signed-rank results at the step level. $\Delta_{t}=(\mathrm{GT}-\mathrm{Hal})_{\mathrm{with}}-(\mathrm{GT}-\mathrm{Hal})_{\mathrm{w/o}}$. NN is the number of paired steps. Holm-adjusted $p$-values are shown for $\Delta_{t}$.
[…]
All tokens 88,706 3.289-3.289 4.826-4.826   0.715\;\;0.715 $10^{-300}$
All tokens (in hallucinated set) 35,116 3.720-3.720 5.273-5.273   0.774\;\;0.774 $1.09{\times}10^{-116}$
Hallucinated steps 190 8.281-8.281 9.435-9.435 0.180-0.180 $0.95$
In the full-token analyses, the median $\Delta$ is positive and highly significant, indicating a systematic shift of the step-level ordering toward GT when vision tokens are present. For the hallucinated-step subset (n=190n{=}190), the median $\Delta$ is small and not significant, i.e., the reversed ordering (Hal >> GT) persists at those positions despite the image.
A.1.4 Figure 4
Setup formulation.
We follow Eq. equation 4 with the final decoder layer (j=Lj{=}L). For each of the 500 MSCOCO validation images, we project every vision token’s final-layer hidden state $h_{i}^{L}$ through the language-modeling head to obtain a distribution over the vocabulary,
$q_{i}(w)\;=\;p_{\theta}(h_{i}^{L},\mathcal{V})\quad(i=1,\dots,|\mathbf{v}|),$
yielding 173,173173{,}173 vision-token projections in total. We evaluate two settings on the same projections: (i) full vocabulary $\mathcal{V}$ and (ii) a CHAIR-derived object subset $\Omega^{(1)}$ of 176 single-token synonyms covering 62 MSCOCO categories denoted as Eq. 9
Main results with confidence intervals.
Table 8 summarizes token-level at-least-one recall@k with 95% Wilson intervals.
Table 8: At-least-one GT recall over N=173,173N{=}173{,}173 vision-token projections with 95% Wilson intervals. GT targets are restricted to the same 62 categories in both settings.
Top-$k$ Full vocabulary Constrained subset (176)
Top-1 2.03% [1.96, 2.10] 40.44% [40.21, 40.67]
Top-5 7.87% [7.74, 8.00] 69.56% [69.34, 69.78]
Top-10 12.38% [12.23, 12.54] 79.78% [79.59, 79.97]
Top-30 21.72% [21.53, 21.91] 89.24% [89.09, 89.39]
A.2 Dataset and evaluation setup
CHAIR evaluation.
CHAIR (Rohrbach et al., 2018) is a generative benchmark designed to measure object hallucinations in image captioning. A hallucinated object refers to any entity mentioned in the generated caption that is not present in the corresponding image. This metric is widely adopted to evaluate the visual grounding capability of vision-language models. We follow the standard CHAIR evaluation procedure based on the MS COCO annotations and the official implementation. [^2] https://github.com/LisaAnne/Hallucination Each image is associated with a set of ground-truth objects defined by the 80 MS COCO detection categories. During evaluation, noun phrases extracted from generated captions are matched against this predefined object list; mentions of out-of-vocabulary objects are ignored, while in-vocabulary objects absent from the ground-truth annotations are considered hallucinated.
Protocol. We use the official CHAIR code and MS COCO annotations, and evaluate on 500 images randomly sampled from the COCO validation set with a fixed seed (seed=42) for reproducibility, prompting each image with “Please describe this image in detail.”.
Metrics. We report the two standard CHAIR metrics and an additional F1:
$\begin{split}\text{CHAIR}_{S}&=\frac{\left|\{\text{hallucinated captions}\}\right|}{\left|\{\text{total captions}\}\right|},\\ \text{CHAIR}_{I}&=\frac{\left|\{\text{hallucinated object mentions}\}\right|}{\left|\{\text{total object mentions}\}\right|},\\ \text{F1}&=\frac{\sum_{\text{images}}\text{F1}_{\text{sample}}}{\left|\{\text{total captions}\}\right|}.\end{split}$
[…]
POPE evaluation.
POPE (Li et al., 2023b) is a discriminative benchmark designed to evaluate the object-level visual grounding capabilities of LVLMs. It formulates the task as a binary object presence question-answering problem, where the model is asked to determine whether a specific object is present in a given image. Evaluation is conducted using standard classification metrics derived from the confusion matrix, including accuracy, precision, and F1 score. Accuracy measures the overall proportion of correct predictions, while precision captures the fraction of predicted positive instances that are indeed correct. F1 score, defined as the harmonic mean of precision and recall, provides a balanced assessment of model performance on this binary classification task. The benchmark covers three datasets: MS COCO (Lin et al., 2014), A-OKVQA (Schwenk et al., 2022), and GQA (Hudson & Manning, 2019). For each dataset, POPE defines three types of query strategies—random, popular, and adversarial—yielding a total of nine evaluation scenarios. Binary questions are constructed in the form “Is there a <object> in the image?”, with the queried object selected according to the designated strategy. Random queries are uniformly sampled from the dataset’s object vocabulary. Popular queries focus on frequently occurring objects, while adversarial queries target objects that are semantically plausible in the given context but are absent from the image, thus probing the model’s reliance on prior co-occurrence statistics rather than visual evidence. Ground-truth labels indicating object presence are provided by the official POPE benchmark. [^3] https://github.com/AoiDragon/POPE For MS COCO, labels are derived from object detection annotations, whereas for A-OKVQA and GQA, they are obtained via SEEM-based segmentation annotations. Each query type consists of 3,000 binary QA instances, evenly split between positive and negative samples. As a result, we evaluate across a total of nine scenarios comprising 27,000 QA examples.
HallusionBench evaluation.
HallusionBench (Guan et al., 2024) is a diagnostic benchmark for image–context reasoning that disentangles Language Hallucination (over-reliance on parametric priors) and Visual Illusion (misinterpretation of visual inputs). All questions are binary (Yes/No) and organized into control pairs to test consistency across visual contexts. We focus on the Visual Dependent (VD) split, which comprises 591 Yes/No questions that require the image to answer; each figure consists of an original image and one or more human-edited variants (e.g., flipping, order reversing, masking, optical-character editing, object editing, color editing) to stress visual robustness. We use the official data and evaluation scripts. [^4] https://github.com/tianyi-lab/HallusionBench
Protocol. Following the official setup, we adopt the text-only GPT-4–assisted judging scheme. Let figures be indexed by $i\in\mathbb{I}$. For each figure $i$, let $\{I_{i,j}\}_{j\in\mathcal{J}_{i}}$ denote its image variants with $j=0$ the original and $j\geq 1$ edited versions, and let $\{q_{i,k}\}_{k\in\mathcal{K}_{i}}$ be the associated Yes/No questions (shared across variants). Given a model’s raw response $\mathcal{M}(I_{i,j},q_{i,k})$, the judge (GPT-4, temperature 0) receives the prompt with $q_{i,k}$ (question), $y_{i,j,k}$ (reference answer), $\mathcal{M}(I_{i,j},q_{i,k})$ (model response)} and returns a label in {Correct, Incorrect, Uncertain}. We binarize correctness as
[…]
Metrics. Let $\mathcal{V}=\{(i,j,k):i\in\mathbb{I},\,j\in\mathcal{J}_{i},\,k\in\mathcal{K}_{i}\}$ be all VD image–question pairs, and let $\mathbb{Q}=\{(i,k):i\in\mathbb{I},\,k\in\mathcal{K}_{i}\}$ be the set of control question pairs. We report the standard HallusionBench accuracies:
$\text{aAcc}\;=\;\frac{1}{|\mathcal{V}|}\sum_{(i,j,k)\in\mathcal{V}}b_{\mathcal{M}}(I_{i,j},q_{i,k}),$
$\text{qAcc}\;=\;\frac{1}{|\mathbb{Q}|}\sum_{(i,k)\in\mathbb{Q}}\mathbf{1}\!\Big\{\,\bigwedge_{j\in\mathcal{J}_{i}}b_{\mathcal{M}}(I_{i,j},q_{i,k})=1\,\Big\},$
$\text{fAcc}\;=\;\frac{1}{\sum_{i\in\mathbb{I}}|\mathcal{J}_{i}|}\sum_{i\in\mathbb{I}}\sum_{j\in\mathcal{J}_{i}}\mathbf{1}\!\Big\{\,\bigwedge_{k\in\mathcal{K}_{i}}b_{\mathcal{M}}(I_{i,j},q_{i,k})=1\,\Big\}.$
Here, aAcc is accuracy per question (overall), qAcc is accuracy per question pair (the same question must be correct across the figure’s image variants), and fAcc is accuracy per image (all questions assigned to a given image—original or edited—must be correct). Following the leaderboard convention, we additionally report Easy aAcc (restricted to original images, j=0j{=}0) and Hard aAcc (restricted to edited images, $j{\geq}1$).
MMMU evaluation.
MMMU (Yue et al., 2024) is a large-scale, multi-discipline VQA benchmark that evaluates college-level multimodal understanding and reasoning across six disciplines, 30 subjects, and heterogeneous image types (charts, diagrams, tables, maps, music sheets, chemical structures, etc.). We use the validation split for controlled comparisons. The validation split includes open-ended and multiple-choice (MC) questions. Since our LLaVA-1.5 configuration supports only single-image inputs, we exclude the 43 multi-image items from the 900 validation questions and report results on the remaining 857 items. Following the official setup, [^5] https://github.com/MMMU-Benchmark/MMMU we append instruction postfixes: for open-ended questions we add “Answer the question using a single word or phrase.”; for MC questions we add “Answer with the option’s letter from the given choices directly.” to adopt official parsing tool.
Metrics. We report overall Accuracy and additionally Subject Win/Tie vs. Greedy, which compares per-subject accuracy against a vanilla greedy-decoding baseline. Let $\mathcal{D}$ be the evaluated set (857 items after filtering) and $y(x)$ the ground-truth answer for item $x\in\mathcal{D}$. For a method $\mathcal{M}$,
$\text{Accuracy}\;=\;\frac{1}{|\mathcal{D}|}\sum_{x\in\mathcal{D}}\mathbf{1}\{\,\mathcal{M}(x)=y(x)\,\}.$
Let $\mathcal{S}$ denote the set of subjects and $\mathcal{D}_{s}\subset\mathcal{D}$ the items belonging to subject $s$. Per-subject accuracy is $\text{Acc}_{\mathcal{M}}(s)=\frac{1}{|\mathcal{D}_{s}|}\sum_{x\in\mathcal{D}_{s}}\mathbf{1}\{\,\mathcal{M}(x)=y(x)\,\}.$ We then count
$\text{Win/Tie}\;=\;\big|\{\,s\in\mathcal{S}:\;\text{Acc}_{\mathcal{M}}(s)\geq\text{Acc}_{\text{Greedy}}(s)\,\}\big|.$
VQAv2 evaluation.
VQAv2 (Goyal et al., 2017) is a large-scale open-ended visual question answering benchmark where each question is annotated with 10 human answers. We use the official evaluation protocol and code. [^6] https://visualqa.org/evaluation.html The metric is designed to be robust to inter-annotator variability by averaging machine accuracy over all $10\choose 9$} subsets of annotators. Concretely, the per-question accuracy assigned to a model answer ans is
$\text{Acc}(\textit{ans})\;=\;\min\!\left\{\frac{\#\text{ of humans that said }\textit{ans}}{3},\,1\right\},$
and the dataset accuracy is the mean of $\text{Acc}(\textit{ans})$ over questions.
LLaVA-Bench-In-the-Wild dataset.
We additionally employ the LLaVA-Bench-In-the-Wild (Liu et al., 2023) dataset to qualitatively evaluate the visual grounding ability of LVLMs in open-ended image understanding. The dataset consists of 24 diverse images, including both real-world photographs and abstract illustrations, and 60 textual prompts designed to elicit complex visual reasoning and language generation. Following prior work Leng et al. (2024), we present representative qualitative examples to illustrate how different decoding strategies affect grounding fidelity and response quality in challenging scenarios.
A.3 Related works
In this section, we provide a detailed discussion of prior works that aim to align visual inputs with language generation. We organize these approaches into three categories—pre-alignment, intra-alignment, and post-alignment—based on the stage where the alignment is applied with respect to the inference process. While this taxonomy follows the temporal flow of inference, the categories are largely orthogonal in methodology and can be composed to enhance multimodal grounding.
A.3.1 Pre-alignment
Recent advances in LVLMs have explored a range of pre-alignment strategies, in which visual and textual modalities are aligned during training to enhance multimodal understanding. Most existing approaches adopt a combination of instruction tuning and architectural refinements to achieve improved visual grounding. For instance, Qwen-VL Bai et al. (2023) and InternVL Chen et al. (2024c) retain decoder-only LLM backbones while incorporating high-resolution vision encoders and cross-modal projection modules. These models leverage large-scale curated image-text data and modality-specific tuning objectives, leading to strong general-purpose performance across a variety of vision-language tasks. Building upon these designs, Qwen2.5-VL Bai et al. (2025) introduces a Dynamic High-Resolution (DHR) processing pipeline and rotary position encodings tailored for vision-language alignment, further enhancing the model’s ability to process diverse image resolutions efficiently. NVLM Dai et al. (2024) explores a broader architectural design space by implementing decoder-only, cross-attention, and hybrid variants, along with a DHR input pipeline and tile tagging mechanisms. On the other hand, there are another line of works designed to mitigate the architectural challenges within the pre-alignment space. Ferret You et al. (2024) focuses on fine-grained spatial referring and grounding by introducing hybrid region representations and a spatial-aware visual sampler, enabling the model to handle points, boxes, and free-form regions. In a different direction, Ovis Lu et al. (2024) addresses the structural mismatch between continuous visual embeddings and discrete textual embeddings. Specifically, Ovis introduces a learnable visual embedding table and a probabilistic vision tokenization scheme that mimics the structure of language token embeddings, thereby aligning the two modalities at the representation level. While these pre-alignment strategies have proven effective, they typically require extensive retraining and are not designed to enable inference-time intervention based on visual context.
A.3.2 Post-alignment
Post-alignment refers to strategies that operate on the generated output sequence, aiming to correct hallucinations or factual inconsistencies after the generation process is complete. Unlike pre- or intra-alignment methods that intervene during training or decoding, post-alignment introduces no changes to the model’s parameters or inference-time logic. A key distinction among post-alignment methods lies in how the correction is applied. Some approaches train an auxiliary revision module, while others rely on external vision models to validate and refine the generated content. For example, LURE Zhou et al. (2023) learns a dedicated revisor that takes the original image and the LVLM-generated caption as input, and outputs a hallucination-corrected caption. The training signal is derived from statistical indicators such as co-occurrence bias and positional uncertainty observed in model outputs. In contrast, Woodpecker Yin et al. (2024) does not require finetuning, but instead performs post-hoc correction through a multi-stage pipeline. Starting from the generated text, it extracts visual concepts, formulates corresponding questions, and verifies them using a pretrained VQA model. Each stage of this process relies on existing pretrained components including question generation, visual verification, and text editing modules to systematically identify and correct hallucinated content. While post-alignment strategies offer a practical way to improve output consistency without altering the base model, they typically rely on additional training or auxiliary models.
A.3.3 Intra-alignment
Improving visual grounding (e.g., mitigating object hallucination) of LVLMs during decoding has evolved through distinct methodological paradigms, each addressing unique aspects of visual-textual misalignment. The research trajectory begins with Contrastive Decoding Li et al. (2023a) strategies, pioneered by VCD Leng et al. (2024), which introduced input distortions to contrast original and perturbed visual distributions, thereby reducing reliance on language priors within LVLMs. This foundation expanded into instruction-aware variants such as ICD Wang et al. (2024), which amplifies alignment uncertainty through textual prompt disturbances, and M3ID Favero et al. (2024), which identifies language priors in the LLM decoder via vision-agnostic input perturbations. Concurrently, attention-centric approaches emerged Gong et al. (2024); Liu et al. (2024b); Woo et al. (2024); Huo et al. (2025); Kang et al. (2025), aiming to mitigate attention misallocation by recalibrating cross-modal attention or introducing contrastive signals for token-level grounding. Other paradigms Chen et al. (2024a); Huang et al. (2024); Wang et al. (2025) have also been explored in recent literature, but all these intra-alignment approaches require additional computation to obtain vision-related reference information.
Algorithm 1 ReVisiT decoding algorithm
Input: LVLM $\mathcal{M}$ with final decoder layer $L$, projection (language modeling) head $W_{\texttt{proj}}$, and vocabulary $\mathcal{V}$), input image II, set of candidate decoder layers $\mathcal{J}$, input text prompt $x_{\text{raw}}$
Tokenize input text: $\mathbf{x}=\{x_{0},\dots,x_{|\mathbf{x}|-1}\}\leftarrow\texttt{Tokenizer}(x_{\text{raw}})$
Encode image: $\mathbf{v}=\{v_{0},\dots,v_{|\mathbf{v}|-1}\}\leftarrow\texttt{VisionEncoder}(I)$
Construct input sequence: $\mathbf{z}=\mathbf{v}\mathbin{\|}\mathbf{x}$ $\triangleright$ $\mathbf{z}=\{z_{0},\dots,z_{T-1}\}$ where $T=|\mathbf{v}|+|\mathbf{x}|$
Initialize output sequence $\mathbf{y}\leftarrow\emptyset$ $\triangleright$ $t\leftarrow 0$
while not end-of-sequence do
  Get decoder hidden states $h_{T+t-1}^{L}$ from $\mathcal{M}$ with context $(\mathbf{z},\mathbf{y})$
  if first-timestep then
   Given $\{h_{i}^{j}\}$ for $i=0,\dots,T-1$ and $j\in\mathcal{J}$ from initial forward pass
   Initialize cache $\mathbf{l}_{\texttt{vision}}\leftarrow\emptyset$
   for each layer $j\in\mathcal{J}$ do
     for $i=0$ to $|\mathbf{v}|-1$ do
      $l_{i}^{j}=W_{\texttt{proj}}^{\top}h_{i}^{j}$
      Append $l_{i}^{j}$ to $\mathbf{l}_{\texttt{vision}}$
     end for
   end for
   Cache matrix $\mathbf{l}_{\texttt{vision}}\in\mathbb{R}^{|\mathcal{J}|\cdot|\mathbf{v}|\times|\mathcal{V}|}$
   first-timestep $\leftarrow$ False
  end if
  Compute vanilla distribution $p_{\text{base}}=\texttt{softmax}_{\mathcal{V}}(W_{\texttt{proj}}^{\top}h_{T+t-1}^{L})$ (Eq. 1)
  Select constrained vocabulary $\mathcal{V}_{\texttt{cons}}^{t}\subseteq\mathcal{V}$ based on APC of $p_{\text{base}}$ (Eq. 3)
  Let $p_{\text{base}}^{\texttt{cons}}=\texttt{softmax}_{\mathcal{V}_{\texttt{cons}}^{t}}(W_{\texttt{proj}}^{\top}h_{T+t-1}^{L})$
  $(j^{*},i^{*})\leftarrow\arg\min_{(j,i)}\texttt{JSD}(p_{\text{base}}^{\texttt{cons}}\|\texttt{softmax}_{\mathcal{V}_{\texttt{cons}}^{t}}(\mathbf{l}_{\texttt{vision}}[j,i]))$
  Let $p_{\text{ref}}^{\texttt{cons}}=\texttt{softmax}_{\mathcal{V}_{\texttt{cons}}^{t}}(\mathbf{l}_{\texttt{vision}}[j^{*},i^{*}])$
  Logit adjustment via element-wise multiplication: $l_{\text{ReVisiT}}=p_{\text{base}}^{\texttt{cons}}\odot p_{\text{ref}}^{\texttt{cons}}$
  Normalize over constrained vocabulary: $p_{\text{ReVisiT}}=\texttt{softmax}_{\mathcal{V}_{\texttt{cons}}^{t}}(l_{\text{ReVisiT}})$
  Sample next token: $y_{t}\sim p_{\text{ReVisiT}}$
  Append $y_{t}$ to $\mathbf{y}$
  $t\leftarrow t+1$
end while
return $\mathbf{y}$
A.4 ReVisiT
In this section, we first provide the pseudo-code of ReVisiT in Algorithm 1, which illustrates how vision tokens are leveraged to refine the output distribution at each decoding step. We highlight that our implementation is memory-efficient, model-agnostic, and requires no additional training or modification to the base model; the decoding function alone is modified. ReVisiT involves two hyperparameters: (1) the pool of decoder layer(s) where vision tokens are selected, and (2) the alignment weight $\alpha$ that determines the size of constrained vocabulary subset. For LLaVA-1.5-7B, we empirically found that using vision tokens from the final decoder layer (“last”) mostly yielded the best grounding fidelity. In contrast, for Qwen2.5-VL-7B, aggregating vision tokens from all even-numbered decoder layers (“all”) was more effective. These selections were fixed per model across all benchmarks. The threshold $\alpha$ controls the size of the constrained vocabulary subset. Lower values (resulting in larger subsets) are used for generative tasks where broader linguistic diversity is desired, while higher values (yielding smaller subsets) are preferred in discriminative tasks to enhance precision. We select $\alpha$ from a logarithmic range between $10^{-1}$ and $10^{-6}$, with $10^{-5}$ serving as the default when task-specific adaptation is not required.
Appendix B Detailed Quantitative Results
B.1 Comprehensive results for 7 models
Table 9: Comprehensive results on CHAIR benchmark across various model sizes and architectures. Lower scores ($\downarrow$) on CHAIRS, CHAIRI and higher ($\uparrow$) F1 indicate better performance. The best results in each model are bolded.
Model Method CHAIRS $\downarrow$ CHAIRI $\downarrow$ F1 $\uparrow$
LLaVA-1.5-7B Greedy 53.8 14.66 82.33
[…]
Qwen2.5-VL-7B Greedy 35.2 8.43 79.85
[…]
InternVL2.5-8B Greedy 33.4 8.47 80.58
[…]
LLaVA-1.5-13B Greedy 48.0 13.29 83.27
[…]
InternVL3-14B Greedy 33.2 9.34 81.49
[…]
InternVL2.5-26B Greedy 31.8 8.39 79.60
[…]
Qwen2.5-VL-32B Greedy 53.6 8.85 82.00
[…]
B.2 Comprehensive Ablation results
Table 10: Ablation study. Panel (a) varies the vision token selection criterion (min/max = min-/max-JSD, random = uniform) and use of a vocabulary subset (full = no constraint, subset = APC). Panel (b) varies the layer scope (last vs all) and the APC threshold $\alpha$. Metrics are CHAIRS (↓), CHAIRI (↓), and F1 (↑). Highlighted rows mark the main configuration.
(a) Vocabulary Constraint $\times$ Token Selection
Vocab Selection CHAIRS $\downarrow$ CHAIRI $\downarrow$ F1 $\uparrow$
subset max 0.0 0.00 0.67
subset random 2.6 7.94 17.63
[…]
full min 0.2 7.14 1.52
(b) Layer Scope $\times$ Threshold $\alpha$
Layer Scope $\bm{\alpha}$ CHAIRS $\downarrow$ CHAIRI $\downarrow$ F1 $\uparrow$
last $1\mathrm{e}{-1}$ 33.8 8.32 80.02
last $1\mathrm{e}{-2}$ 37.0 9.13 79.88
last $1\mathrm{e}{-3}$ 34.0 7.88 80.97
last $1\mathrm{e}{-4}$ 34.2 8.21 80.41
last $1\mathrm{e}{-5}$ 31.4 8.04 80.12
last $1\mathrm{e}{-6}$ 32.8 7.21 79.68
all $1\mathrm{e}{-1}$ 33.8 8.82 80.70
all $1\mathrm{e}{-2}$ 36.8 8.91 79.93
all $1\mathrm{e}{-3}$ 38.8 9.36 79.60
all $1\mathrm{e}{-4}$ 32.8 8.17 80.17
[…]
all $1\mathrm{e}{-6}$ 32.4 9.38 79.90
B.3 Comprehensive POPE results
Table 11: Comprehensive results on POPE benchmark across benchmarks (MS-COCO, A-OKVQA, GQA) and setups (Random / Popular / Adversarial). Higher scores ($\uparrow$) on Accuracy, Precision, and F1 indicate better performance.
Setup Method LLaVA-1.5-7B Qwen2.5-VL-7B
Acc $\uparrow$ Prec $\uparrow$ F1 $\uparrow$ Acc $\uparrow$ Prec $\uparrow$ F1 $\uparrow$
MS-COCO Random Greedy 89.07 89.54 89.00 84.07 99.71 81.09
DoLa 89.00 89.58 88.92 81.53 99.89 77.37
VCD 87.23 86.62 87.34 84.57 99.15 81.88
M3ID 89.07 89.54 89.00 83.93 99.32 80.96
CODE 86.53 87.23 86.41 84.00 98.02 81.26
SID 89.67 91.49 89.43 86.93 99.03 85.10
[…]
MS-COCO Popular Greedy 85.63 83.72 86.03 83.47 97.99 80.52
DoLa 85.63 83.85 86.00 81.10 98.54 76.96
VCD 83.53 80.75 84.25 83.93 97.39 81.27
M3ID 85.63 83.72 86.03 83.40 97.81 80.46
CODE 84.73 84.14 84.86 83.63 97.02 80.92
SID 86.13 85.19 86.32 86.23 97.22 84.42
[…]
MS-COCO Adversarial Greedy 79.27 74.72 81.01 83.10 97.06 80.16
DoLa 79.40 74.97 81.08 80.83 97.73 76.71
VCD 77.67 72.90 79.77 83.37 95.96 80.73
M3ID 79.27 74.72 81.01 83.00 96.79 80.06
CODE 78.67 75.23 80.02 83.00 95.41 80.31
SID 80.30 76.56 81.59 85.17 94.74 83.39
[…]
A-OKVQA Random Greedy 86.33 81.00 87.42 87.80 96.86 86.49
DoLa 86.43 81.25 87.47 85.33 96.99 83.26
VCD 83.70 77.91 85.23 87.87 96.56 86.62
M3ID 86.33 81.00 87.42 87.40 95.91 86.11
CODE 83.80 78.48 85.18 87.50 93.64 86.55
SID 87.57 83.48 88.28 88.83 96.12 87.88
[…]
A-OKVQA Popular Greedy 78.77 71.74 81.72 85.93 92.58 84.74
DoLa 79.00 72.06 81.85 83.37 92.17 81.43
VCD 76.47 69.58 79.99 85.90 92.10 84.78
M3ID 78.77 71.74 81.72 85.77 92.21 84.59
CODE 77.70 71.17 80.68 86.07 90.62 85.24
SID 80.63 74.30 82.87 87.00 92.11 86.16
[…]
A-OKVQA Adversarial Greedy 68.20 61.86 74.91 80.83 82.59 80.30
DoLa 68.33 62.00 74.95 79.63 84.22 78.17
VCD 67.80 61.69 74.47 81.10 82.69 80.63
M3ID 68.20 61.86 74.91 79.90 81.00 79.54
CODE 68.30 62.23 74.61 79.37 78.73 79.59
SID 71.33 64.75 76.57 82.40 83.38 82.14
[…]
GQA Random Greedy 85.97 79.72 87.30 87.53 97.15 86.12
DoLa 86.07 79.92 87.36 83.47 94.66 81.10
VCD 83.77 77.16 85.53 87.83 96.78 86.55
M3ID 86.00 79.74 87.33 87.43 96.14 86.12
CODE 84.87 78.58 86.37 87.07 94.48 85.89
SID 85.97 79.72 87.30 88.80 96.78 87.88
[…]
GQA Popular Greedy 73.80 66.38 78.64 83.80 88.82 82.68
DoLa 73.97 66.56 78.73 78.57 83.71 76.80
VCD 71.87 64.76 77.32 83.33 87.09 82.44
M3ID 73.83 66.39 78.67 80.67 82.39 80.14
CODE 73.90 66.60 78.60 81.47 83.29 80.95
SID 73.80 66.38 78.64 84.27 87.25 83.61
[…]
GQA Adversarial Greedy 68.17 61.60 75.19 81.50 84.36 80.70
DoLa 68.40 61.80 75.30 78.37 83.32 76.63
VCD 67.63 61.22 74.82 82.00 84.63 81.29
M3ID 68.20 61.62 75.22 80.50 82.11 80.00
CODE 68.70 62.12 75.39 79.87 80.56 79.64
SID 68.17 61.60 75.19 82.37 83.79 81.99
[…]
B.4 Token-Level computation speed
Table 12: Per-token inference latency (ms/token) of different decoding strategies. We report the mean ± standard deviation over 300 samples.
Method LLaVA-1.5-7B Qwen2.5-VL-7B
Greedy 26.0 ± 0.2 72.3 ± 1.5
DoLa 34.2 ± 0.4 82.6 ± 0.4
VCD 51.7 ± 0.4 143.4 ± 1.1
M3ID 49.5 ± 0.2 143.3 ± 13.2
CODE 81.0 ± 6.3 176.2 ± 45.1
SID 51.9 ± 0.4 146.5 ± 1.3
[…]
Appendix C Detailed Qualitative Results
C.1 Case studies for vision token analysis
To support the qualitative visualization presented in Figures 5 and 2, we provide the detailed numerical values and positional metadata used in our vision token analysis. All examples are drawn from the LLaVA-Bench-In-the-Wild dataset, using the LLaVA-1.5-7B. Vision token embeddings are extracted from decoder layer L=32L={32}, and the threshold for constrained vocabulary selection is set to $\alpha=10^{-5}$.
First, Figure 5 presents a case study from the LLaVA-Bench-In-the-Wild dataset (image ID: 002.jpg, query: “How many uncut fruits are in the image?”). At decoding step $t=2$, ReVisiT selects vision token index 229 as the most relevant reference based on the constrained vocabulary distribution. Table 13 reports the log-probability values from three sources: (1) the base model output (Eq. 1), (2) the vision token projection at index 229 (Eq. 4), and (3) the final combined distribution after log-probability fusion (Eq. 6). These values are computed over the constrained subset $\mathcal{V}_{\texttt{cons}}^{2}$. The token ‘‘three’’ is ultimately selected, as the vision token significantly boosts its relevance, increasing its probability from 35.45% to 58.27%, while suppressing ‘‘four’’ from 50.78% to 38.21%.
[…]
The token on the right (vision token index 193) is selected to demonstrate the actual decoding behavior of ReVisiT. At generation step $t=72$, this token serves acts as a reference informer to refine the output distribution through log-probability fusion. The final token ‘‘painting’’ is selected, with its probability increasing from 8.08% (base) to 39.15% (final), while alternatives such as hallucinatory token ‘‘person’’ are suppressed. Detailed values are presented in Table 15.
Table 13: Detailed numerical values for Figure 5. Token scores over the constrained vocabulary $\mathcal{V}_{\texttt{cons}}^{2}$ at decoding step $t=2$ with ReVisiT are presented. The selected vision token (index 229) amplifies the relevance of “three”, increasing its probability from 35.45% (base) to 58.27% (final), while suppressing “four” from 50.78% to 38.21%.
Rank Token Base Logit (229) Vision Token Logit Final Logit
Log-Prob Prob (%) Log-Prob Prob (%) Log-Prob Prob (%)
1 three 1.04-1.04 $35.45$ 1.47-1.47 $22.98$ 2.51-2.51 $58.27$
2 four 0.68-0.68 $50.78$ 2.25-2.25 $10.52$ 2.93-2.93 $38.21$
3 two 4.57-4.57 $1.04$ 1.19-1.19 $30.44$ 5.76-5.76 $2.26$
4 five 2.54-2.54 $7.91$ 4.45-4.45 $1.16$ 6.99-6.99 $0.66$
5 a 4.33-4.33 $1.31$ 3.08-3.08 $4.58$ 7.42-7.42 $0.43$
6 six 3.94-3.94 $1.94$ 5.15-5.15 $0.58$ 9.10-9.10 $0.08$
7 5.10-5.10 $0.61$ 4.22-4.22 $1.47$ 9.32-9.32 $0.06$
8 several 5.75-5.75 $0.32$ 5.39-5.39 $0.45$ 11.15-11.15 $0.01$
9 seven 5.79-5.79 $0.31$ 6.12-6.12 $0.22$ 11.90-11.90 $0.00$
10 un 8.65-8.65 $0.02$ 3.80-3.80 $2.24$ 12.45-12.45 $0.00$
11 at 8.28-8.28 $0.03$ 4.77-4.77 $0.85$ 13.05-13.05 $0.00$
12 one 10.05-10.05 $0.00$ 3.69-3.69 $2.50$ 13.74-13.74 $0.00$
13 in 11.15-11.15 $0.00$ 2.99-2.99 $5.01$ 14.14-14.14 $0.00$
14 eight 7.38-7.38 $0.06$ 6.78-6.78 $0.11$ 14.16-14.16 $0.00$
15 f 9.52-9.52 $0.01$ 4.76-4.76 $0.86$ 14.28-14.28 $0.00$
16 multiple 7.66-7.66 $0.05$ 6.66-6.66 $0.13$ 14.32-14.32 $0.00$
17 only 8.68-8.68 $0.02$ 5.89-5.89 $0.28$ 14.57-14.57 $0.00$
18 nine 7.71-7.71 $0.04$ 6.97-6.97 $0.09$ 14.68-14.68 $0.00$
19 half 8.62-8.62 $0.02$ 6.22-6.22 $0.20$ 14.85-14.85 $0.00$
20 many 8.70-8.70 $0.02$ 6.21-6.21 $0.20$ 14.91-14.91 $0.00$
21 all 10.38-10.38 $0.00$ 4.85-4.85 $0.78$ 15.23-15.23 $0.00$
22 still 11.11-11.11 $0.00$ 4.15-4.15 $1.57$ 15.26-15.26 $0.00$
23 an 10.33-10.33 $0.00$ 5.10-5.10 $0.61$ 15.43-15.43 $0.00$
24 some 10.05-10.05 $0.00$ 5.40-5.40 $0.45$ 15.46-15.46 $0.00$
25 more 11.15-11.15 $0.00$ 4.68-4.68 $0.92$ 15.83-15.83 $0.00$
26 about 10.05-10.05 $0.00$ 5.90-5.90 $0.27$ 15.95-15.95 $0.00$
27 both 10.74-10.74 $0.00$ 5.23-5.23 $0.53$ 15.98-15.98 $0.00$
28 the 11.98-11.98 $0.00$ 4.07-4.07 $1.71$ 16.05-16.05 $0.00$
29 not 11.58-11.58 $0.00$ 4.59-4.59 $1.02$ 16.16-16.16 $0.00$
30 over 11.99-11.99 $0.00$ 4.25-4.25 $1.42$ 16.25-16.25 $0.00$
31 different 11.93-11.93 $0.00$ 4.73-4.73 $0.88$ 16.66-16.66 $0.00$
32 ten 8.56-8.56 $0.02$ 8.15-8.15 $0.03$ 16.71-16.71 $0.00$
33 just 10.97-10.97 $0.00$ 5.78-5.78 $0.31$ 16.75-16.75 $0.00$
34 fruit 11.81-11.81 $0.00$ 5.04-5.04 $0.65$ 16.85-16.85 $0.00$
35 as 12.06-12.06 $0.00$ 4.82-4.82 $0.81$ 16.88-16.88 $0.00$
36 small 12.11-12.11 $0.00$ 4.86-4.86 $0.78$ 16.97-16.97 $0.00$
37 various 10.39-10.39 $0.00$ 6.71-6.71 $0.12$ 17.10-17.10 $0.00$
38 also 11.94-11.94 $0.00$ 5.21-5.21 $0.55$ 17.15-17.15 $0.00$
39 very 11.80-11.80 $0.00$ 5.54-5.54 $0.39$ 17.35-17.35 $0.00$
40 around 11.28-11.28 $0.00$ 6.07-6.07 $0.23$ 17.35-17.35 $0.00$
41 numerous 10.02-10.02 $0.00$ 7.51-7.51 $0.05$ 17.54-17.54 $0.00$
42 no 10.70-10.70 $0.00$ 6.93-6.93 $0.10$ 17.64-17.64 $0.00$
43 total 11.17-11.17 $0.00$ 6.80-6.80 $0.11$ 17.97-17.97 $0.00$
44 currently 11.17-11.17 $0.00$ 6.89-6.89 $0.10$ 18.06-18.06 $0.00$
45 cut 11.78-11.78 $0.00$ 6.46-6.46 $0.16$ 18.24-18.24 $0.00$
46 quite 11.62-11.62 $0.00$ 6.71-6.71 $0.12$ 18.33-18.33 $0.00$
47 few 11.91-11.91 $0.00$ 6.64-6.64 $0.13$ 18.55-18.55 $0.00$
48 twelve 10.07-10.07 $0.00$ 8.53-8.53 $0.02$ 18.60-18.60 $0.00$
49 lots 11.52-11.52 $0.00$ 7.37-7.37 $0.06$ 18.89-18.89 $0.00$
50 th 11.34-11.34 $0.00$ 7.63-7.63 $0.05$ 18.98-18.98 $0.00$
51 actually 11.84-11.84 $0.00$ 7.40-7.40 $0.06$ 19.24-19.24 $0.00$
52 approximately 11.05-11.05 $0.00$ 8.45-8.45 $0.02$ 19.50-19.50 $0.00$
53 exactly 12.14-12.14 $0.00$ 7.48-7.48 $0.06$ 19.62-19.62 $0.00$
54 plenty 11.78-11.78 $0.00$ 7.95-7.95 $0.04$ 19.73-19.73 $0.00$
55 eleven 10.82-10.82 $0.00$ 9.25-9.25 $0.01$ 20.07-20.07 $0.00$
Table 14: Detailed numerical values for Figure 2 (left). Log-probabilities and probabilities (%) of selected constrained vocabulary tokens as projected by three vision tokens (pink, yellow, cyan) under the full vocabulary $\mathcal{V}$ and the constrained vocabulary $\mathcal{V}_{\texttt{env}}$.
Vision Token Top (Pink) Middle (Yellow) Bottom (Cyan)
Vocab $\mathcal{V}$ $\mathcal{V}_{\texttt{env}}$ $\mathcal{V}$ $\mathcal{V}_{\texttt{env}}$ $\mathcal{V}$ $\mathcal{V}_{\texttt{env}}$
Token Log-Prob % Log-Prob % Log-Prob % Log-Prob % Log-Prob % Log-Prob %
[…]
mountain 14.07-14.07 $0.00$ 5.48-5.48 $0.42$ 6.47-6.47 $0.15$ 3.22-3.22 $3.99$ 15.02-15.02 $0.00$ 5.46-5.46 $0.43$
person 13.39-13.39 $0.00$ 4.80-4.80 $0.82$ 12.27-12.27 $0.00$ 9.01-9.01 $9.86 0.010.01 \cellcolorgray$$0.30 \cellcolorgray!20$$73.81$
[…]
tree 12.06-12.06 $0.00$ 3.48-3.48 $3.09$ 11.71-11.71 $0.00$ 8.46-8.46 $0.02$ 12.47-12.47 $0.00$ 2.92-2.92 $5.42$
cloud 13.88-13.88 $0.00$ 5.30-5.30 $0.50$ 9.52-9.52 $0.01$ 6.26-6.26 $0.19$ 12.46-12.46 $0.00$ 2.91-2.91 $5.47$
sea 13.80-13.80 $0.00$ 5.21-5.21 $0.55$ 6.82-6.82 $0.11$ 3.57-3.57 $2.81$ 13.24-13.24 $0.00$ 3.68-3.68 $2.52$
grass 15.41-15.41 $0.00$ 6.82-6.82 $0.11$ 7.19-7.19 $0.08$ 3.94-3.94 $1.95$ 13.41-13.41 $0.00$ 3.86-3.86 $2.11$
rock 13.35-13.35 $0.00$ 4.77-4.77 $0.85$ 4.95-4.95 $0.71$ 1.70-1.70 $18.33$ 12.41-12.41 $0.00$ 2.85-2.85 $5.77$
Table 15: Detailed numerical values for Figure 2 (right). Token scores over the constrained vocabulary $\mathcal{V}_{\texttt{cons}}^{72}$ at decoding step $t=72$ with ReVisiT. The selected vision token (index $193$) amplifies the relevance of “painting”, increasing its probability from 8.08% (base) to 39.15% (final) while suppressing hallucinatory token “person” from 15.10% to 7.19%.
Rank Token Base Logit (193) Vision Token Logit Final Logit
Log-Prob Prob (%) Log-Prob Prob (%) Log-Prob Prob (%)
1 painting 2.52-2.52 $8.08$ 1.27-1.27 $27.99$ 3.79-3.79 $39.15$
2 rock 2.60-2.60 $7.42$ 2.28-2.28 $10.22$ 4.88-4.88 $13.11$
3 mountain 2.14-2.14 $11.76$ 3.03-3.03 $4.83$ 5.17-5.17 $9.82$
4 person 1.89-1.89 $15.10$ 3.59-3.59 $2.75$ 5.48-5.48 $7.19$
5 landscape 2.27-2.27 $10.30$ 3.42-3.42 $3.27$ 5.70-5.70 $5.82$
6 d 3.45-3.45 $3.16$ 2.29-2.29 $10.14$ 5.74-5.74 $5.55$
7 p 3.91-3.91 $2.01$ 2.45-2.45 $8.60$ 6.36-6.36 $3.00$
8 scene 3.80-3.80 $2.23$ 2.70-2.70 $6.75$ 6.50-6.50 $2.60$
9 small 2.83-2.83 $5.91$ 3.70-3.70 $2.48$ 6.52-6.52 $2.54$
10 beautiful 3.67-3.67 $2.54$ 3.03-3.03 $4.84$ 6.70-6.70 $2.13$
11 large 2.93-2.93 $5.34$ 3.78-3.78 $2.28$ 6.71-6.71 $2.11$
12 scen 3.57-3.57 $2.81$ 3.17-3.17 $4.19$ 6.74-6.74 $2.04$
13 hill 4.01-4.01 $1.82$ 2.83-2.83 $5.91$ 6.84-6.84 $1.86$
14 chair 2.85-2.85 $5.78$ 4.38-4.38 $1.26$ 7.23-7.23 $1.26$
15 bow 3.87-3.87 $2.09$ 3.68-3.68 $2.53$ 7.54-7.54 $0.92$
16 second 3.80-3.80 $2.23$ 4.52-4.52 $1.09$ 8.33-8.33 $0.42$
17 smaller 3.02-3.02 $4.90$ 6.04-6.04 $0.24$ 9.06-9.06 $0.20$
18 boat 3.00-3.00 $4.98$ 6.34-6.34 $0.18$ 9.34-9.34 $0.15$
19 distant 4.14-4.14 $1.59$ 5.43-5.43 $0.44$ 9.57-9.57 $0.12$
C.2 Without vocabulary subset case study
Figure 8: Without vocabulary subset case study. Qualitative case study from Qwen2.5-VL-7B. w/o subset refers to ablation result of without vocabulary subset constraint, whereas w/ subset refers to our proposed ReVisiT.
C.3 Qualitative examples
To illustrate the qualitative improvements, we present examples from the LLaVA-Bench-In-The-Wild dataset, which consists of diverse real-world images requiring complex reasoning beyond conventional benchmarks. As shown in Figure 9, the input image is a cartoon illustration requiring contextual and conceptual understanding. While baseline methods such as vanilla greedy decoding fail to capture the main theme and generate overly generic or literal descriptions, or even getting stuck in repetitive phrasing. In contrast, ReVisiT successfully identifies the underlying concept: about neural network layers.
Figure 9: Qualitative example. The input image is a cartoon-style illustration contrasting classical statistical learning and neural network reasoning via a visual metaphor, emphasizing the shift from theoretical rigor to the heuristic of “stacking more layers.” We compare the generated responses of vanilla greedy decoding, M3ID, and ReVisiT, highlighting how ReVisiT better captures the intended visual analogy compared to the baseline methods.
Figure 10: Additional qualitative example. The input image is a illustration showing a bear, a cat, and a rabbit seated around a table with a plate of donuts. We compare the responses of vanilla greedy decoding and ReVisiT to the question, “What are the animals in the painting and what are they doing?” While the greedy output introduces a hallucinated detail (“cookie”) and assigns actions not visually supported (e.g., “the cat drinking from a cup”), ReVisiT provides a more faithful description aligned with the visual content, accurately identifying the animals and their activities.
Appendix D Broader Impacts and Limitations
Broader impacts.
Like most approaches built upon LLMs, the proposed method inherits the limitations and potential societal risks of the underlying models, including biases in training data, reinforcement of stereotypes, and the generation of factually incorrect or harmful content. While our decoding strategy improves visual grounding by constraining generation based on vision token signals, it does not explicitly eliminate such biases and may inadvertently propagate them through vision-conditioned outputs. On the other hand, by incorporating grounded visual evidence during generation, our method may contribute to mitigating hallucinations and increasing the interpretability of multimodal outputs. We believe that this is an important step toward trustworthy and reliable deployment of LVLMs in the real-world.
Limitations.
A potential limitation of our method is its strong reliance on the semantics encoded in vision tokens, which may lead to overfitting to visually salient cues. While this behavior improves visual grounding in most cases, it can result in an overemphasis on perceptually grounded elements at the expense of implicit, commonsense, or non-visual contextual information. To mitigate such overfitting issue, future work could explore adaptive interpolation strategies between reference logits and vanilla logits, or incorporate auxiliary confidence calibration mechanisms to balance multimodal importance dynamically. Another limitation of our approach is its exclusive reliance on the model’s internal representations without access to external knowledge sources. As a result, it cannot correct factual errors or hallucinations originating from pretraining data, particularly in cases where visual evidence alone is insufficient for disambiguation or grounding. While our method improves the alignment between visual inputs and generated text, it does not incorporate retrieval mechanisms that could supplement the model with up-to-date or context-specific information. Nevertheless, due to its modular and decoding-time nature, our method is compatible with retrieval-augmented generation (RAG) pipelines and could be integrated with external knowledge modules in future work.